OpenStudy (isaiahlobo):

Solve and check. C/3 = 6 1/7 A.2 1/21 B.2 1/7 C.18 1/7 D.18 3/7

jaynator495 (jaynator495):

i beleive you meant @Jacob902 isaih

Elsa213 (elsa213):

What do you think the correct answer is?

OpenStudy (isaiahlobo):

@Jacob902

OpenStudy (isaiahlobo):

i think A

OpenStudy (isaiahlobo):

lol

Elsa213 (elsa213):

Why do you think the correct answer is A?

OpenStudy (isaiahlobo):

divide

OpenStudy (mathmale):

Please show all of your work, so that the rest of us could give you meaningful guidance.

jaynator495 (jaynator495):

C/3 = 6 1/7 Im assuming this means the answer you choose devided by 3 = 6 1/7?

OpenStudy (isaiahlobo):

i dont know how to do this tha much

jaynator495 (jaynator495):

If that is what you mean, then unfortantly A is not the correct answer...

jaynator495 (jaynator495):

What divided by 3 is 6?

OpenStudy (mathmale):

Solve and check. C/3 = 6 1/7. Exactly what are you supposed to find, and what are the steps of the process?

jaynator495 (jaynator495):

Mathmale i can help them on this if they mean what i think they mean ._.

OpenStudy (mrnood):

\[\frac{ C }{ 3 } = 6\frac{ 1 }{ 7 } \] so \[\frac{ C }{ 3} = (6 + \frac{ 1 }{ 7 })\] Multiply both sides of the equation by 3 to get the answer

OpenStudy (isaiahlobo):

D. then

jaynator495 (jaynator495):

Correct! :)

OpenStudy (isaiahlobo):

yaa

OpenStudy (isaiahlobo):

thx